ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

UCC Ghana and UCA Arkansas 1991 MS Classification: not given Asiedu-Addo, Samuel K., University College of Education, Winneba, Ghana, Griffith, Linda K., Mullins, Charles W., Wiredu, Emmanuel, Bonsu, Osei Mensan Perceptions and attitudes related to mathematics learning: A comparison between UCC Ghana and UCA Arkansas The question addressed in this study is “Is there a difference in the perceptions and attitudes about mathematics learning between students at the University of Cape Coast in Ghana, Africa and students at the University of Central Arkansas in Conway, Arkansas?” The researchers selected and developed survey instruments that were completed by students at both universities. Differences do exist between UCA and UCC mathematics students’ perceptions and attitudes about learning mathematics. These are most pronounced for students taking less rigorous college mathematics courses. The attitudes of the UCC students are more positive and the UCA students are less anxious about mathematics.